# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ V Conclusion and Perspectives
In this work, we explore different quantum-feature kernel implementations (1) and we study their performances on two benchmark datasets using a support vector machine (SVM). In a first step we introduce a graph-embedding scheme that not only allows for the edge features to be embedded in a neutral-atom Rydberg Hamiltonian (2), but also for the node features (the atomic masses) via an appropriate choice of the local detuning fields, cf. Subsec. II.2. Hence, we study the effect of node-feature embedding [time evolution via $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10)] versus non-node-feature embedding [time evolution via $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9)]. Having demonstrated theoretically (cf. proposition 1) and empirically (cf. Fig. 4) that node-feature embedding enhances the kernel‘s expressiveness, we observe a comparable SVM-prediction performance between the two cases (cf. Fig. 3).
In a second step, we compare two conceptually different quantum-feature kernels: the quantum evolution kernel (QEK) $\kappa_{\mathrm{QEK}}$ (12) [13, 14] which operates with a global observable, the excitation probability $P_{k}^{\mathcal{G}}$ (6), and the proposed generalized-distance quantum-correlation (GDQC) kernel $\kappa_{\mathrm{GDQC}}$ (13) which draws from the correlation matrix $C_{i,j}^{\mathcal{G}}$ (6), a local observable. While these two kernels could perform qualitatively different for different graph-related prediction tasks, for the global prediction task studied in this work, both kernels lead to a comparable performance which is on par with standard classical state-of-the-art algorithms (cf. Tab. 1).
In a third step, we attempt to further enhance the quantum-feature kernels‘ performance. For this purpose we combine the information of observables sampled at multiple steps across the time evolution by means of pooling operations (cf. Subsec. III.3). Here, we find that pooling can significantly enhance the models performance enabling the quantum-feature kernels to surpass the classical baselines (cf. Tabs. 1-2). Together, the presented quantum-feature kernel implementations enlarge the scope of the previously existing techniques in the field of quantum graph machine learning.
Hyper-parameter sensitivity remains an open question. We notice that it is difficult to draw definitive conclusions on the global tendencies of our models depending on parameter choices, such as the local detuning value $\delta_{0}$ [cf. Eq. (8)], the Rabi frequency $\Omega$ [cf. Eq. (2)], as well as the time steps at which to sample the observables (regardless whether they are pooled or not). This motivates the need of automatic, data-driven strategies to set these hyper-parameters such as gradient-based optimization, Bayesian- or evolutionary search.
@@ Acknowledgements.
The author gratefully acknowledges Louis-Paul Henry, Vittorio Vitale, and Francesco Musso for the many fruitful discussions that have significantly contributed to the development of this work. We would also like to thank Anton Quelle, Kemal Bidzhiev, Mauro Mendizabal, and Elie Merhej for their valuable contributions to the development of the emulation tools, which were essential for the implementation of the algorithms presented in this work.
Appendix A Improved graph discrimination using local detuning
@@
Appendix B Properties of GDQC
In this appendix we prove the proposition 2: The kernel function $\kappa_{\mathrm{GDQC}}$ defined Eq. (13) is: (1) a kernel function (symmetric positive semi-definite), (2) valid on graph data for any graph distance $d_{\mathcal{G}}$, and (3) equal to 1 on any isomorphic pair $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
Let us recall that the generalized-distance quantum-correlation kernel (GDQC) $\kappa_{\mathrm{GDQC}}$ (13) is defined via a scalar product
@@ Appendix B Properties of GDQC
where $C^{\mathcal{G}}\left(t\right)$ denotes the correlation matrix (6) associated with the graph $\mathcal{G}$, cf. Eqs. (1), (9), (10), and $D^{\mathcal{G}}$ is the generalized-distance matrix. The index iterates over the whole dataset of $N_{\mathrm{data}}$ graphs, i.e. $\mu,\mu^{\prime}\in\left\{1,2,\dots,N_{\mathrm{data}}\right\}$. Naturally, the correlation matrix $C^{\mathcal{G}}\left(t\right)$ is a real-valued observable since it is derived from a Hermitian operator. Consequently, each element of the $N_{\chi}$-dimensional GDQC feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ is real-valued with $N_{\chi}=N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D}$. By continuation, the GDQC kernel $\kappa_{\mathrm{GDQC}}\left(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}}\right)$ is also real-valued. Next, we prove the 3 statements of proposition 2 point by point.
@@
confirming that the GDQC kernel is indeed positive semi-definite.
(2). We demonstrate that the kernel is invariant under node label permutations. Hereby, it suffices to show that the GDQC feature vector satisfies
@@
Consequently, the permutation of a graph is identical to the inverse permutation of the matrix indices associated with the non-permuted graph. Such a transformation, however, has no effect on the GDQC feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$, as the feature-vector elements are simply counts of the number of values falling into a predefined interval, cf. algorithm 1 and Fig. 2. Thus, the GDQC kernel $\kappa_{\mathrm{GDQC}}$ (22) is invariant under node label permutations. It remains to prove that the transformation relations (25) are actually true.
@@
for the kernel function (22). The property (34) follows trivially from the GDQC feature vector invariance (24), and the normalization of the feature vector $\left\|\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]\right\|=1$. ∎
Appendix C Expressiveness of GDQC
In this Appendix we prove the proposition 3: The kernel function $\kappa_{\mathrm{GDQC}}$ (13) is as expressive than the GD-WL test given the latter is initialized with colors obtained from the multi-set of the values of the correlation matrix rows/columns.

### v2

@@ V Conclusion and Perspectives
In this work, we investigate different quantum-feature kernel implementations (1) and evaluate their performance on two benchmark datasets using support vector machines. First, we introduce a graph-embedding scheme that encodes edge information into a neutral-atom Rydberg Hamiltonian (2), and additionally incorporates node features (atomic masses) via an appropriate choice of local detuning fields (cf. Subsec. II.2). This allows us to compare node-feature embedding (time evolution under $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$, Eq. (10)) with a non-attributed construction (time evolution under $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$, Eq. (9)). While node-feature embedding increases kernel expressiveness both theoretically (Prop. 1) and empirically (Fig. 4), we observe comparable SVM prediction performance in the two settings (Fig. 3), with a slight advantage for the local addressability approach.
Second, we compare two conceptually different quantum-feature kernels: the quantum evolution kernel $\kappa_{\mathrm{QEK}}$ (Eq. (15)) [24, 2], which is based on the global excitation-probability observable $P_{k}^{\mathcal{G}}(t)$ (Eq. (6)), and the proposed generalized-distance quantum-correlation kernel $\kappa_{\mathrm{GDQC}}$ (Eq. (17)), which is based on the local correlation matrix $C_{i,j}^{\mathcal{G}}(t)$ (Eq. (6)). While these kernels can, in principle, behave differently depending on the target task, for the global classification tasks considered here they achieve comparable performance and are competitive with strong classical baselines (Tab. 1).
Third, we improve performance by pooling information from observables sampled at multiple times across the evolution (cf. Subsec. III.4). Pooling can significantly enhance model performance, enabling the quantum-feature kernels to surpass the classical baselines (Tabs. 1-2). Overall, the proposed implementations broaden the scope of existing approaches in quantum graph machine learning.
Our quantum kernels do not provide an asymptotic speedup over the classical baselines considered here: $\mathrm{QEK}$ scales as $\mathcal{O}(N_{\mathrm{shots}}\,N)$, while $\mathrm{GDQC}$ requires $\mathcal{O}(N_{\mathrm{shots}}\,N^{2})$. By comparison, the best-performing classical baseline in our experiments is the WL optimal assignment kernel, with complexity $\mathcal{O}(h\,M+N)$. Moreover, the cost of analog quantum evolution is difficult to quantify in standard algorithmic terms, since complexity notions are inherently tied to digital computation. In this work, the value proposition therefore lies in empirical performance (competitive weighted F1 scores on both benchmarks) rather than an asymptotic runtime advantage.
Finally, hyper-parameter sensitivity remains an important open question. In particular, we find it difficult to draw definitive conclusions about global performance trends as a function of parameters such as the local detuning scale $\delta_{0}$ (Eq. (8)), the Rabi frequency $\Omega$ (Eq. (2)), the $\mathrm{GDQC}$ binning resolution (Alg. 1), and the choice of sampling times (with or without pooling).
@@ Acknowledgements.
The authors gratefully acknowledge Louis-Paul Henry, Vittorio Vitale, and Francesco Musso for the many fruitful discussions that have significantly contributed to the development of this work. We also thank Anton Quelle, Kemal Bidzhiev, Mauro Mendizabal, and Elie Merhej for their valuable contributions to the development of the emulation tools, which were essential for the implementation of the algorithms presented in this work.
Appendix A Time-dependent perturbation theory
In this appendix, we revisit basics on the time-dependent perturbation theory, and we derive the leading observable contributions which are presented in Subsec. II.4.
We begin with a generic system Hamiltonian
$\hat{\mathcal{H}}\left(t\right)$ $=\hat{H}_{0}+\hat{V}\left(t\right),$ (25)
decomposed into an unperturbed part $\hat{H}_{0}$ (assumed to be time independent, for simplicity), and a time-dependent perturbation part $\hat{V}\left(t\right)$ which is considered to be small. In this part we distinguish between the Schrödinger and the interaction picture, which we label via subscripts $\mathrm{S}$ and $\mathrm{I}$, respectively. In the Schrödinger picture, the time-evolved state is given by
$\left|\psi_{\mathrm{S}}\left(t\right)\right\rangle$ $=\hat{U}_{\mathrm{S}}\left(t,t_{0}\right)\left|\psi_{\mathrm{S}}\left(t_{0}\right)\right\rangle,$ (26)
with the Schrödinger time evolution operator
$\hat{U}_{\mathrm{S}}\left(t,t_{0}\right)$ $=\hat{\mathcal{T}}_{t}\,\exp\left[-\frac{\mathsf{i}}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\;\hat{\mathcal{H}}\left(t^{\prime}\right)\right]$ (27)
$=e^{-\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t}\hat{U}_{\mathrm{I}}\left(t,t_{0}\right)e^{\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t_{0}},$ (28)
whose relation to the interaction-picture time evolution operator $\hat{U}_{\mathrm{I}}\left(t,t_{0}\right)$ is shown in Eq. (28). The Schrödinger equation in the interaction picture becomes
$\mathsf{i}\hbar\frac{\partial}{\partial t}\hat{U}_{\mathrm{I}}\left(t,t_{0}\right)$ $=\hat{V}_{\mathrm{I}}\left(t\right)\hat{U}_{\mathrm{I}}\left(t,t_{0}\right),$ (29)
with
$\hat{V}_{\mathrm{I}}\left(t\right)$ $=e^{\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t}\hat{V}\left(t\right)e^{-\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t},$ (30)
and the formal solution
$\hat{U}_{\mathrm{I}}\left(t,t_{0}\right)$ $=\hat{\mathcal{T}}_{t}\;\exp\left[-\frac{\mathsf{i}}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\,\hat{V}_{\mathrm{I}}\left(t^{\prime}\right)\right],$ (31)
where $\hat{\mathcal{T}}_{t}$ is the time ordering operator. At this point we assume that the exponent in (31) is effectively small such that the expansion
$\hat{U}_{\mathrm{I}}\left(t,t_{0}\right)$ $=\sum_{l=0}^{\infty}\hat{U}_{\mathrm{I}}^{\left(l\right)}\left(t,t_{0}\right),$ (32)
in terms of a Neumann series is justified. The $l$-th series coefficient is given by
$\hat{U}_{\mathrm{I}}^{\left(l\right)}\left(t,t_{0}\right)$ $=\left(-\frac{\mathsf{i}}{\hbar}\right)^{l}\int_{t_{0}}^{t}dt_{1}\int_{t_{0}}^{t_{1}}dt_{2}\cdots\int_{t_{0}}^{t_{l-1}}dt_{l}\,\times$
$\qquad\hat{V}_{\mathrm{I}}\left(t_{1}\right)\hat{V}_{\mathrm{I}}\left(t_{2}\right)\cdots\hat{V}_{\mathrm{I}}\left(t_{l}\right),$ (33)
where $\hat{U}_{\mathrm{I}}^{\left(0\right)}\left(t,t_{0}\right)=\mathbbm{1}$. For practical purposes, it is often convenient to exploit the inductive relation
$\hat{U}_{\mathrm{I}}^{\left(l\right)}\left(t,t_{0}\right)$ $=-\frac{\mathsf{i}}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\;\hat{V}_{\mathrm{I}}\left(t^{\prime}\right)\;\hat{U}_{\mathrm{I}}^{\left(l-1\right)}\left(t^{\prime},t_{0}\right),$ (34)
for $l\geq 1$. Switching back to the Schrödinger picture via relation (28) we can finally expand the time-evolved wave function (26) according to
$\left|\psi_{\mathrm{S}}\left(t\right)\right\rangle$ $=\sum_{l=0}^{\infty}\left|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\right\rangle,$ (35)
where the $l$-th order wave-function contribution is given by
$\left|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\right\rangle$ $=e^{-\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t}\hat{U}_{\mathrm{I}}^{\left(l\right)}\left(t,t_{0}\right)e^{\frac{\mathsf{i}}{\hbar}\hat{H}_{0}t_{0}}\left|\psi_{\mathrm{S}}\left(t_{0}\right)\right\rangle.$ (36)
In practice, the computation can be facilitated using the induction rule (34) such that
$\left|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\right\rangle$ $=-\frac{\mathsf{i}}{\hbar}\int_{t_{0}}^{t}dt^{\prime}\;e^{-\frac{\mathsf{i}}{\hbar}\hat{H}_{0}\left(t-t^{\prime}\right)}\hat{V}\left(t^{\prime}\right)\,\left|\psi_{\mathrm{S}}^{\left(l-1\right)}\left(t^{\prime}\right)\right\rangle,$ (37)
for $l\geq 1$ and $|\psi_{\mathrm{S}}^{\left(0\right)}\left(t\right)\rangle=\exp\left[-\frac{\mathsf{i}}{\hbar}\hat{H}_{0}\left(t-t_{0}\right)\right]\left|\psi_{\mathrm{S}}\left(t_{0}\right)\right\rangle$. Now, wave-function contributions and eventually, observables can be evaluated order by order. Typically, the evaluation of (37) is conducted using the eigenbasis of the unperturbed Hamiltonian $\hat{H}_{0}$.
Starting from the Rydberg Hamiltonian (2) we consider the transverse part as the perturbation, i.e.
$\hat{H}_{0}$ $=\frac{1}{2}\sum\nolimits_{i,j}^{\prime}J_{ij}\hat{n}_{i}\hat{n}_{j}-\hbar\sum_{i=1}^{N}\delta_{i}\hat{n}_{i},$ $\hat{V}$ $=\frac{\hbar\Omega}{2}\sum_{i=1}^{N}\hat{\sigma}_{i}^{x}.$ (38)
Note that $\hat{V}$ is time independent in this case. The goal is to derive the leading contributions in powers of $\left(\Omega t\right)$ for the observables of interest (6), the excitation probability $P_{k}\left(t\right)$, the site occupation $n_{i}\left(t\right)$, and the correlation matrix $C_{i,j}\left(t\right)$. Recall that our initial state at time $t_{0}=0$ is the all-zero state, i.e. $\left|\psi\left(t_{0}=0\right)\right\rangle=\left|\bm{0}\right\rangle$. By construction, the $l$-th order wave-function term $|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\rangle$ scales as $\mathcal{O}\left((\Omega t)^{l}\right)$, see Eq. (37). Furthermore, because the perturbation $\hat{V}$ (38) can maximally cause one excitation per application, the contribution $|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\rangle$ can maximally contain $l$ excitations. For later convenience, we explicitly evaluate the first three wave-function contributions with $l\in\left\{0,1,2\right\}$. Using the relation (37), one finds $|\psi_{\mathrm{S}}^{\left(0\right)}\left(t\right)\rangle=\left|\bm{0}\right\rangle$, and
$\left|\psi_{\mathrm{S}}^{\left(1\right)}\left(t\right)\right\rangle$ $=\frac{\Omega}{2}\sum_{i=1}^{N}\frac{1-e^{\mathsf{i}\delta_{i}t}}{\delta_{i}}\,\hat{\sigma}_{i}^{x}\left|\bm{0}\right\rangle,$ (39)
$\left|\psi_{\mathrm{S}}^{\left(2\right)}\left(t\right)\right\rangle$ $=a_{0}^{\left(2\right)}\,\left|\bm{0}\right\rangle+\sum\nolimits_{i,j}^{\prime}a_{ij}^{\left(2\right)}\,\hat{\sigma}_{j}^{x}\,\hat{\sigma}_{i}^{x}\,\left|\bm{0}\right\rangle,$ (40)
where we defined the second-order amplitudes
$a_{0}^{\left(2\right)}$ $=\frac{\Omega^{2}}{4}\sum_{i=1}^{N}\,\frac{e^{\mathsf{i}\delta_{i}t}-1-\mathsf{i}\delta_{i}t}{\delta_{i}^{2}},$ (41)
$a_{ij}^{\left(2\right)}$ $=-\frac{\Omega^{2}}{4\delta_{i}}\left[\frac{1-e^{-\mathsf{i}t\omega_{ij}^{\left(2\right)}}}{\omega_{ij}^{\left(2\right)}}-\frac{e^{\mathsf{i}\delta_{i}t}-e^{-\mathsf{i}\omega_{ij}^{\left(2\right)}t}}{\omega_{ij}^{\left(2\right)}+\delta_{i}}\right],$ (42)
with $\omega_{ij}^{\left(2\right)}=(J_{ij}/\hbar)-\delta_{i}-\delta_{j}$.
Next, we evaluate the observables, starting with the excitation probability operator $\hat{P}_{k}=\sum_{\bm{b}}\delta_{k,\sum_{i}b_{i}}\left|\bm{b}\right\rangle\left\langle\bm{b}\right|$, cf. Eq. (6). Since the index $k$ effectively counts the number of excitations, it is clear that the first non-vanishing projection occurs for the wave-function contribution $|\psi_{\mathrm{S}}^{\left(l=k\right)}\left(t\right)\rangle$, leading to
$P_{k}\left(t\right)$ $=\sum_{l,l^{\prime}=k}^{\infty}\Big\langle\psi_{\mathrm{S}}^{(l^{\prime})}\left(t\right)\Big|\hat{P}_{k}\Big|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\Big\rangle=\mathcal{O}\left(\left(\Omega t\right)^{2k}\right).$ (43)
In this work, we only include contributions up to order $\mathcal{O}\left(\left(\Omega t\right)^{2}\right)$. Hence, the two non-negligible terms are $P_{0}\left(t\right)$ and $P_{1}\left(t\right)$ which can be unpacked as
$P_{0}\left(t\right)$ $=1+2\Re\left[\left\langle\bm{0}\big|\psi_{\mathrm{S}}^{\left(2\right)}\left(t\right)\right\rangle\right]+\mathcal{O}\left(\left(\Omega t\right)^{4}\right),$ (44)
$P_{1}\left(t\right)$ $=\sum_{i=1}^{N}\left|\left\langle\bm{0}\,\big|\,\hat{\sigma}_{i}^{x}\,\big|\,\psi_{\mathrm{S}}^{\left(1\right)}\left(t\right)\right\rangle\right|^{2}+\mathcal{O}\left(\left(\Omega t\right)^{4}\right).$ (45)
Here, we have exploited that $|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\rangle$ contains an even (odd) number of excitations for even (odd) order $l$. After insertion of Eqs. (39)-(40), the expressions (44)-(45) simplify to the results (12)-(13) shown in the main text. Next, we consider the site occupation operator $\hat{n}_{i}$ whose action on a state $\left|0_{i}\right\rangle$ is $\hat{n}_{i}\left|0_{i}\right\rangle=0$. Correspondingly, the order $l=0$ does not contribute to $n_{i}\left(t\right)$ and one finds
$n_{i}\left(t\right)$ $=\sum_{l,l^{\prime}=1}^{\infty}\Big\langle\psi_{\mathrm{S}}^{(l^{\prime})}\left(t\right)\Big|\hat{n}_{i}\Big|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\Big\rangle=\mathcal{O}\left(\left(\Omega t\right)^{2}\right)$
$=\left\langle\psi_{\mathrm{S}}^{\left(1\right)}\left(t\right)\right|\hat{n}_{i}\left|\psi_{\mathrm{S}}^{\left(1\right)}\left(t\right)\right\rangle+\mathcal{O}\left(\left(\Omega t\right)^{4}\right),$ (46)
with the evaluated expression shown in Eq. (14). Lastly, we compute the correlation matrix $C_{i,j}\left(t\right)$. Surely, its diagonal elements $C_{i,i}\left(t\right)=n_{i}\left(t\right)$ are described by Eq. (46). Since $\hat{n}_{i}\hat{n}_{j}\,|\psi_{\mathrm{S}}^{\left(l=\left\{0,1\right\}\right)}\left(t\right)\rangle=0$, for the off-diagonal elements one obtains
$C_{i,j\neq i}\left(t\right)$ $=\sum_{l,l^{\prime}=2}^{\infty}\Big\langle\psi_{\mathrm{S}}^{(l^{\prime})}\left(t\right)\Big|\hat{n}_{i}\hat{n}_{j}\Big|\psi_{\mathrm{S}}^{\left(l\right)}\left(t\right)\Big\rangle=\mathcal{O}\left(\left(\Omega t\right)^{4}\right),$ (47)
i.e. for the employed expansion order, the off-diagonal elements $C_{i,j}\left(t\right)$ can be neglected.
Appendix B Classical graph kernels
To deal with the classical kernel methods introduced in Subsec. III.1, we employ external implementations for standard graph kernels, namely the Python library GraKeL. Here, we report the parameter and performance choices that are made to produce the benchmarks listed in table 1.
For all kernels, we choose to normalize the kernel matrix. This improves numerical stability and makes the subsequent optimization less sensitive to an arbitrary scaling of the kernel. In GraKeL, this corresponds to the normalize option, which applies the following transformation:
$K_{ij}^{\text{norm}}=K_{ij}\big/\sqrt{K_{ii}K_{jj}},$ (48)
where in principle, it should hold $K_{ii}>0$. In practice, there can be special cases where $K_{ii}=0$ or $K_{ii}<0$ due to anomalies in the dataset or to numerical instabilities. In such a case, we regularize the kernel matrix before normalizing via $K\rightarrow K+\epsilon\mathbbm{1}$ with $\epsilon\ll 1$. Such a case occurs for the Graphlet Subsampling kernel on the PTC_FM dataset where we choose $\epsilon=10^{-3}$ for regularization.
For the Random Walk kernel, we use the exponential variant rather than the geometric one. The geometric formulation relies on the inversion of a matrix $(\mathbbm{1}-\lambda A_{\times})$, whose numerical stability is highly sensitive to the hyperparameter $\lambda$. For this reason, we employ the exponential formulation which, although computationally more expensive (complexity $\mathcal{O}(N^{6})$ versus $\mathcal{O}(N^{3})$ for the geometric variant), is substantially more stable in practice.
Unless stated otherwise, all hyperparameters are left to their default values. We only report non-default settings below.
  • Random Walk (kernel_type=exponential):
    • MUTAG: lamda=5
    • PTC_FM: lamda=10.
  • Graphlet Sampling (normalize=False):
    • MUTAG: $k=9$
    • PTC_FM: $k=7$.
  • ShortestPath (algorithm_type=floyd_warshall).
Appendix C Improved graph discrimination using local detuning
@@
Appendix D Properties of $\mathrm{GDQC}$
In this appendix we prove the proposition 2: The function $\kappa_{\mathrm{GDQC}}$ defined in Eq. (17) is a valid graph kernel, i.e. it is symmetric positive semi-definite on the space of graphs equipped with any graph distance $d_{\mathcal{G}}$, and is isomorphism-invariant such that $\kappa_{\mathrm{GDQC}}(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}})=1$ whenever $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$ are isomorphic. Let us recall that the generalized-distance quantum-correlation kernel $\kappa_{\mathrm{GDQC}}$ (17) is defined via a scalar product
@@ Appendix D Properties of $\mathrm{GDQC}$
where $C^{\mathcal{G}}\left(t\right)$ denotes the correlation matrix (6) associated with the graph $\mathcal{G}$, cf. Eqs. (1), (9), (10), and $D^{\mathcal{G}}$ is the generalized-distance matrix. The index iterates over the whole dataset of $N_{\mathrm{data}}$ graphs, i.e. $\mu,\mu^{\prime}\in\left\{1,2,\dots,N_{\mathrm{data}}\right\}$. Naturally, the correlation matrix $C^{\mathcal{G}}\left(t\right)$ is a real-valued observable since it is derived from a Hermitian operator. Consequently, each element of the $N_{\chi}$-dimensional $\mathrm{GDQC}$ feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$ is real-valued with $N_{\chi}=N_{\mathrm{bins}}^{C}N_{\mathrm{bins}}^{D}$. By continuation, the $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}\left(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}}\right)$ is also real-valued. Next, we prove the 3 statements of proposition 2 point by point.
@@
confirming that the $\mathrm{GDQC}$ kernel is indeed positive semi-definite.
(2). We demonstrate that the kernel is invariant under node label permutations. Hereby, it suffices to show that the $\mathrm{GDQC}$ feature vector satisfies
@@
Consequently, the permutation of a graph is identical to the inverse permutation of the matrix indices associated with the non-permuted graph. Such a transformation, however, has no effect on the $\mathrm{GDQC}$ feature vector $\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]$, as the feature-vector elements are simply counts of the number of values falling into a predefined interval, cf. algorithm 1 and Fig. 2. Thus, the $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}$ (52) is invariant under node label permutations. It remains to prove that the transformation relations (55) are actually true.
@@
for the kernel function (52). The property (64) follows trivially from the $\mathrm{GDQC}$ feature vector invariance (54), and the normalization of the feature vector $\left\|\bm{\chi}\left[C^{\mathcal{G}}\left(t\right),D^{\mathcal{G}}\right]\right\|=1$. ∎
Appendix E Expressiveness of $\mathrm{GDQC}$
In this Appendix we prove the proposition 3: The $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}$ matches the expressiveness of the GD-WL refinement [54] when GD-WL is initialized with node colors constructed from the multisets of (binned) correlation values in the rows/columns of $C^{\mathcal{G}}(t)$.